Solve for x: 3^x = 4
Select an answer below
a.log 4/3
b.log 4/3
c.log 4-log 3
d.ln 4/ln 3
Solve for x: 9^x+2 = 27
Select an answer below
a.1
b.2/3
c.-1/2
d.-2

Answer by edjones(8007)

a^x=y
log(a)y=x
3^x = 4
log(3)4=x
.
9^x+2 = 27
9^x=25
x=log(9)25
=log(10)25/log(10)9
=1.46497
Ed